OpenStudy (anonymous):

Solve this system of equations using the addition method. x+y=3 2x-y=0 @erikalove951

OpenStudy (dan815):

they want you to add these 2 equations up

OpenStudy (dan815):

x+y = 3 2x-y=0 so adding x+y to 2x-y = adding 3 to 0

OpenStudy (anonymous):

but i dont get them still

OpenStudy (dan815):